OpenStudy (anonymous):

Solve Trigonometric equation:- sinx +cosx +sin2x=1/2.............reply fast....

OpenStudy (anonymous):

sin2x can be rewritten

OpenStudy (anonymous):

ya as 2sinxcosx but what next?

OpenStudy (shubhamsrg):

2sinx + 2cos x + 4sinx cosx = 1 2sinx( 1 + 2cosx) + 2cosx = 1 2sinx(1+ 2cosx) + (1+2 cosx) =2 (1+ 2sinx)(1+2cosx) = 2

OpenStudy (shubhamsrg):

I don't have much idea next, leme see

OpenStudy (anonymous):

yes thanks a lot thats absolutely fine! i got an idea to solve the question

OpenStudy (shubhamsrg):

Will you share that idea ?

OpenStudy (anonymous):

2sinx=1; 2cosx=1 and further by the formulae sin theta =sin alpha and so on....

OpenStudy (shubhamsrg):

sinx and cosx both can not be 1/2 for the same value of x, you're logic is flawed.

OpenStudy (shubhamsrg):

Its quite complicated to solve this, are you sure you have the right ques ?

OpenStudy (anonymous):

i know there may be some problem in the question or the expansion....wait..question is cosx+sinx-sin2x=1/2 so in that way it can be solved

OpenStudy (shubhamsrg):

-sin2x !!

OpenStudy (shubhamsrg):

its easily solvable then

OpenStudy (shubhamsrg):

2cosx + 2sinx - 4sinx cosx = 1 2sinx(1 - 2cosx) = (1-2cosx) =>(2sinx)(1- 2cosx) =0 BAZINGA !

OpenStudy (anonymous):

wow!!! thanks....i was a bit confused....thanks a lot!

OpenStudy (shubhamsrg):

Glad to have helped.

OpenStudy (shubhamsrg):

Correction :

OpenStudy (shubhamsrg):

it'll be (1-2sinx)(1- 2cosx) =0

OpenStudy (anonymous):

ya i got that....k thanks bye gtg